### v1

# Can Large Language Models Improve Spectral Graph Neural Networks?

Abstract
Spectral Graph Neural Networks (SGNNs) have attracted significant attention due to their ability to approximate arbitrary filters. They typically rely on supervision from downstream tasks to adaptively learn appropriate filters. However, under label-scarce conditions, SGNNs may learn suboptimal filters, leading to degraded performance. Meanwhile, the remarkable success of Large Language Models (LLMs) has inspired growing interest in exploring their potential within the GNN domain. This naturally raises an important question: Can LLMs help overcome the limitations of SGNNs and enhance their performance? In this paper, we propose a novel approach that leverages LLMs to estimate the homophily of a given graph. The estimated homophily is then used to adaptively guide the design of polynomial spectral filters, thereby improving the expressiveness and adaptability of SGNNs across diverse graph structures. Specifically, we introduce a lightweight pipeline in which the LLM generates homophily-aware priors, which are injected into the filter coefficients to better align with the underlying graph topology. Extensive experiments on benchmark datasets demonstrate that our LLM-driven SGNN framework consistently outperforms existing baselines under both homophilic and heterophilic settings, with minimal computational and monetary overhead.
1 Introduction
The past decade has witnessed great success and prosperity of graph neural networks (GNNs) in diverse data science and engineering scenarios, such as traffic network (Hu et al. 2019; Jiang and Luo 2022), abnormal detection (Tang et al. 2022, 2024a), relational databases (Cappuzzo, Papotti, and Thirumuruganathan 2020; Huang et al. 2022) and recommender systems (Sharma et al. 2024; Gao et al. 2023). Existing GNNs can be broadly categorized into spatial GNNs and spectral GNNs. Spatial GNNs often adopt a message-passing mechanism to learn node representations by aggregating neighbor node features. In contrast, spectral GNNs map node features to a new desired space by selectively attenuating or amplifying the Fourier coefficients induced by the normalized Laplacian matrix. This study primarily centers on the realm of spectral GNNs.
In real-world scenarios, graphs often exhibit varying degrees of homophily. In homophilic graphs, nodes of the same class are more likely to be connected, whereas in heterophilic graphs, connections tend to form between nodes of different classes. Prior studies [24, 37, 42] have shown that heterophilous structures pose significant challenges for spatial GNNs based on message passing, as these models typically rely on the implicit assumption of homophily. To address this issue, many methods have been proposed from the spatial perspective [40]. In contrast, spectral GNNs [3, 19] with learnable filters offer a promising alternative. By learning spectral filters directly from the graph structure, these models can better adapt to heterophilic settings and mitigate the aggregation of noisy information from dissimilar neighbors.
However, since the labels of real-world graph datasets may be very sparse, existing spectral graph neural network methods will learn inappropriate filters, resulting in suboptimal performance. As shown in Figure 1, the optimal filter on the Cora dataset is a low-pass filter due to its high homophily. However, classical spectral graph neural networks such as GPRGNN (Chien et al. 2021), BernNet (He et al. 2021), and JacobiConv (Wang and Zhang 2022) learn high-pass filters. Therefore, the performance of spectral graph neural networks may be suboptimal.
In recent years, large language models (LLMs), exemplified by GPT-4 (Achiam et al. 2023), have shown impressive proficiency in comprehending and reasoning over textual information, thereby revolutionizing a wide range of domains, including natural language processing (Zhao et al. 2023), computer vision (Wu et al. 2023), and graph representation learning (Ren et al. 2024)]. In this work, we explore the potential of language models in spectral graph neural networks. To the best of our knowledge, this is the first study to leverage large language models to improve spectral graph neural networks. Specifically, we aim to address the following research questions:
1. Can large language models (LLMs) enhance spectral GNNs? LLMs are well known for their prowess in understanding natural language, while spectral GNNs excel at node classification in heterophilic graphs. Effectively combining the text comprehension capabilities of LLMs with the filtering power of spectral GNNs presents a significant challenge. Existing efforts to integrate GNNs and LLMs typically take one of two approaches: either feeding textual graphs directly into LLMs—limiting the model’s ability to leverage graph structure, or fine-tuning LLMs for graph-specific tasks, which introduces substantial computational and deployment costs.
2. Can LLMs effectively guide and be integrated into existing spectral GNNs? Spectral GNNs have the powerful ability to approximate arbitrary filters, which are particularly suitable for both homophilic and heterophilic graphs. However, these models usually rely on supervision from downstream tasks to learn appropriate spectral filters, which can lead to suboptimal filter learning in label-scarce or weakly-supervised settings. On the other hand, LLMs demonstrate strong performance in language understanding and text-driven tasks, as seen in models like LLM-GNN (Chen et al. 2023b). This opens up a promising and novel direction: using LLM-generated predictions to assist and improve spectral GNNs.
In this paper, we propose an LLM-based enhancement framework for spectral GNNs. Specifically, to address the first research question: spectral GNNs may learn inappropriate filters due to sparse labels, while the homophily ratio serves as a global characteristic of a graph dataset. We naturally leverage LLMs to predict the homophily ratio, enabling spectral GNNs to perceive global graph properties even under limited supervision. Notably, predicting homophily only requires sampling around 100 edges, making the cost per dataset as low as \$1, without the need for costly fine-tuning or deployment.
For the second research question, we incorporate the predicted homophily ratio into existing polynomial spectral GNNs to enhance their performance. Specifically, we construct various heterophily-aware bases according to the predicted homophily, and then integrate them into polynomial spectral GNNs. This integration allows the model to fully utilize homophily information, boosting overall performance.
Our contributions are summarized as follows:
  • To the best of our knowledge, this is the first work to investigate the integration of LLMs into spectral GNNs. We show that LLMs can effectively predict the homophily ratio to guide spectral GNNs, thereby boosting their performance without the need for fine-tuning or additional deployment overhead.
  • We introduce a novel LLM-based enhancement framework for spectral GNNs, which is compatible with various LLMs and polynomial spectral architectures. By leveraging the global semantic understanding of LLMs, our method significantly improves the adaptability and performance of spectral GNNs across diverse scenarios.
  • Extensive experiments on multiple benchmark datasets demonstrate that our method consistently improves the performance of existing spectral GNNs, while incurring minimal computational and financial overhead.
2 Related Work
This section introduces related research, including spectral GNNs and LLM for Graph.
Spectral GNNs. According to whether the filter can be learned, the spectral GNNs can be divided into pre-defined filters and learnable filters. In the category of pre-defined filters, GCN (Kipf and Welling 2017) uses a simplified first-order Chebyshev polynomial. APPNP (Gasteiger, Bojchevski, and Günnemann 2019) utilizes Personalized Page Rank to set the weight of the filter. In the category of learnable filters. ChebNet (Defferrard, Bresson, and Vandergheynst 2016) uses Chebyshev polynomials with learnable coefficients. GPR-GNN (Chien et al. 2021) extends APPNP by directly parameterizing its weights. BernNet (He et al. 2021) uses Bernstein polynomials to learn filters and forces all coefficients positive. JacobiConv (Wang and Zhang 2022) adopts an orthogonal and flexible Jacobi basis to accommodate a wide range of weight functions. ChebNetII (He, Wei, and Wen 2022) uses Chebyshev interpolation to learn filters. Specformer (Bo et al. 2023) performs self-attention in the spectral domain to learn a set-to-set spectral filter. Recently, some works have applied spectral GNNs to node-level filtering. For example, DSF (Guo et al. 2023) proposes a novel diversified spectral filtering framework that automatically learns node-specific filter weights. UniFilter (Huang et al. 2024) integrates heterophilic and homophilic bases to construct a universal polynomial basis, UniBasis, which partially alleviates the problems of over-smoothing and over-squashing.
LLM for Graph. Existing research on applying large language models (LLMs) to graph learning can be broadly categorized into GNN-centric and LLM-centric approaches. GNN-centric methods leverage LLMs to extract node features from raw data, and then use GNNs for downstream prediction tasks (He et al. 2024; Xie et al. 2023). In contrast, LLM-centric methods integrate GNNs to enhance the performance of LLMs in graph-related tasks (Tang et al. 2024b; Zhang et al. 2024a). Some studies also employ LLMs to assign edge weights in text-attributed graphs (Sun et al. 2023; Ling et al. 2024). However, these methods are not specifically designed for heterophilic graphs. Beyond these, a few studies aim to enhance GNNs by leveraging LLMs to identify meaningful or noisy edges. For instance, GraphEdit (Guo et al. 2024) utilizes LLMs for graph structure learning by detecting and removing noisy connections; LLM4RGNN (Zhang et al. 2024b) identifies malicious and critical edges to improve the adversarial robustness of GNNs; and LLM4HeG (Wu et al. 2024) integrates LLMs into GNNs for heterophilic TAGs through edge discrimination and adaptive edge reweighting.
In contrast to existing methods, the proposed method uniquely focuses on leveraging LLMs to improve and enhance spectral GNNs, which has been underexplored in prior work.
3 Preliminaries
3.1 Spectral GNN
Assume we are given an undirected homogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{n}\right\}$ denotes the set of $n$ nodes, $\mathcal{E}$ is the edge set, and $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the node feature matrix. The corresponding adjacency matrix is $\mathbf{A}\in\{0,1\}^{n\times n}$, where $\mathbf{A}_{ij}=1$ if there is an edge between nodes $v_{i}$ and $v_{j}$, and $\mathbf{A}_{ij}=0$ otherwise. The degree matrix $\mathbf{D}=\text{diag}(d_{1},\ldots,d_{n})$ is a diagonal matrix where the $i$-th diagonal entry is $d_{i}=\sum_{j}\mathbf{A}_{ij}$. The normalized Laplacian matrix is defined as $\mathbf{\hat{L}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, where $\mathbf{I}$ denotes the identity matrix. The normalized adjacency matrix is $\mathbf{\hat{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. Let $\mathbf{\hat{L}}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ denote the eigen-decomposition of $\mathbf{\hat{L}}$, where $\mathbf{U}$ is the matrix of eigenvectors and $\boldsymbol{\Lambda}=\text{diag}([\lambda_{1},\lambda_{2},\ldots,\lambda_{n}])$ is the diagonal matrix of eigenvalues.
Spectral GNNs are based on the Fourier transform in signal processing. The Fourier transform of a graph signal $\mathbf{x}$ is given by $\hat{\mathbf{x}}=\mathbf{U}^{\top}\mathbf{x}$, and its inverse is $\mathbf{x}=\mathbf{U}\hat{\mathbf{x}}$. Accordingly, the graph convolution of the signal $\mathbf{x}$ with a kernel $\mathbf{g}$ can be defined as:
$\mathbf{z}=\mathbf{g}*_{\mathcal{G}}\mathbf{x}=\mathbf{U}\left((\mathbf{U}^{\top}\mathbf{g})\odot(\mathbf{U}^{\top}\mathbf{x})\right)=\mathbf{U}\mathbf{\hat{G}}\mathbf{U}^{\top}\mathbf{x},$ (1)
where $\mathbf{\hat{G}}=\operatorname{diag}(\hat{g}_{1},\ldots,\hat{g}_{n})$ denotes the spectral kernel coefficients. To avoid explicit eigen-decomposition, recent works approximate different kernels $\mathbf{H}$ using polynomial functions $h(\cdot)$ as follows:
$\mathbf{H}=h(\mathbf{\hat{L}})=h_{0}\mathbf{\hat{L}}^{0}+h_{1}\mathbf{\hat{L}}^{1}+h_{2}\mathbf{\hat{L}}^{2}+\cdots+h_{K}\mathbf{\hat{L}}^{K},$ (2)
where $K$ is the order of the polynomial $h(\cdot)$ and $h_{K}$ is the coefficient of the $K$-th order term. Thus, Eq. (1) can be rewritten as:
$\mathbf{Z}=\mathbf{H}\mathbf{X}=h(\mathbf{\hat{L}})\mathbf{X}=\mathbf{U}h(\boldsymbol{\Lambda})\mathbf{U}^{\top}\mathbf{X},$ (3)
where $\mathbf{Z}$ is the output (prediction) matrix. According to Eq. (2), Eq. (3), and recent studies (Chen et al. 2023a), a $K$-th order polynomial in spectral GNNs is equivalent to aggregating information from $K$-hop neighbors in spatial GNNs.
3.2 Homophily
The homophily metric measures the degree of association between connected nodes. Homophily can be measured in various ways, such as node-level homophily, edge-level homophily, and class-level homophily. The widely adopted edge homophily (Zhu et al. 2020) is defined as follows:
$\mathcal{H}_{\text{edge }}(\mathcal{G})=\frac{1}{|\mathcal{E}|}\sum_{(u,v)\in\mathcal{E}}\mathbf{1}\left(y_{u}=y_{v}\right),$ (4)
where $\mathbf{1}(\cdot)$ is the indicator function, i.e., $\mathbf{1}(\cdot)=1$ if the condition holds, otherwise $\mathbf{1}(\cdot)=0$. $y_{u}$ is the label of node $u$, and $y_{v}$ is the label of node $v$ $|\mathcal{E}|$ is the size of the edge set.
4 Methodology
Figure 1: The overall framework of the proposed model. First, the proposed method samples a subset of edges (node pairs) from the dataset. Then we use LLM to predict whether the connected nodes belong to the same class. Based on the LLM’s predictions, we estimate the graph’s homophily. Finally, a heterophily basis is constructed using the predicted homophily and integrated into existing polynomial spectral GNNs to perform the node classification task.
This section describes the proposed method, which consists of two modules: (1) using an LLM to predict the homophily, and (2) incorporating the predicted homophily into existing spectral GNNs. The following provides a detailed introduction to these two modules.
4.1 Estimating Homophily Using LLM
It is well established that, for a graph, the homophily ratio measures the degree of correlation among neighboring nodes and reflects the proportion of edges connecting nodes of the same class. Consequently, accurately estimating the homophily ratio is critical for better leveraging the graph’s structural properties and enabling targeted spectral filtering. For example, in a highly homophilic graph, a polynomial spectral GNN should ideally learn a low-pass filter.
Table 1: Comparison of homophily predicted by the LLM with varying prompts, with indicators for whether CoT and Most-voting are used.
Prompt Type Cora Citeseer Pubmed History
COT Most-voting
× × 0.41 0.63 0.72 0.58
× 0.51 0.76 0.67 0.65
× 0.37 0.63 0.81 0.55
0.70 0.81 0.83 0.77
Ground Truth 0.81 0.78 0.80 0.66
For simplicity, we adopt the widely used edge-level homophily as our estimation target, as Eq. (4). Specifically, we aim to estimate the proportion of homophily edges (i.e., edges linking nodes of the same class) among all edges. However, it is impractical to evaluate the homophily of every edge, especially in large-scale graphs with millions of edges, due to the high inference cost associated with LLMs. Therefore, we propose to estimate the overall homophily ratio by sampling a subset of edges.
To achieve accurate estimation of the homophily ratio, it is important to assess the reliability of predictions, such as through calibrated confidence scores. Motivated by recent advances (Chen et al. 2023b) in generating calibrated confidence from LLMs, we investigate the following strategies:
  • Vanilla prompting: Directly querying the model for its confidence.
  • Reasoning-based prompting(CoT): Guiding the model through annotation generation using techniques such as chain-of-thought and multi-step reasoning.
  • Consistency-based prompting (Most-voting): Querying the model multiple times and selecting the most frequent prediction via majority voting.
  • Hybrid prompting: Combining Reasoning-based and consistency-based methods to enhance robustness.
As shown in Table 1, both reasoning-based prompting and consistency-based prompting contribute to improved performance in predicting homophily. Therefore, we adopt a hybrid prompting strategy that integrates their respective strengths to further enhance the accuracy of homophily estimation. For example, the prompt used for the Cora dataset is designed as follows, while prompts for other datasets can be found in Appendix A.
[…]
After obtaining the homophily labels for a subset of edges using reasoning-based prompting, we further apply consistency-based prompting by querying each sample five times. If at least three out of five predictions yield the same answer, we consider it the final prediction for that sample. The process is illustrated as follows:
$y_{e}=\left\{\begin{array}[]{ll}1&\text{ if }r_{e}\geq 3\\ 0&\text{ if }r_{e}<3\end{array},\right.$
Once the predictions for each sampled edge are obtained, we compute the predicted edge-level homophily $\hat{h}$ as follows:
$\hat{h}=\frac{|\{(u,v)\in\mathcal{E}_{s}\mid\text{LLM}(u,v)=\text{True}\}|}{|\mathcal{E}_{s}|},$ (5)
where $\mathcal{E}_{s}$ denotes the set of all sampled edges.
4.2 Combination of Homophily with Existing Polynomial Filters
After predicting the homophily ratio using an LLM, we incorporate the estimated value into an existing polynomial spectral GNN. To enable spectral GNNs to leverage the homophily information predicted by Eq. (5), we construct a set of heterophily-aware basis vectors inspired by UniFilter (Huang et al. 2024). Specifically, the angle between each pair of basis vectors is defined as follows:
$\theta=\frac{\pi}{2}(1-\hat{h}).$
Thus, the angle between different heterophily-aware basis vectors $\mathbf{u}_{i}$ and $\mathbf{u}_{j}$ is given by:
$\mathbf{u}_{i}\cdot\mathbf{u}_{j}=\begin{cases}\cos\theta=\cos\left(\frac{(1-\hat{h})\pi}{2}\right)&\text{ if }i\neq j,\\ 1&\text{ if }i=j.\end{cases}$
After obtaining the heterophily basis vectors set $\left\{\mathbf{u}_{0},\mathbf{u}_{1},\cdots,\mathbf{u}_{K}\right\}$, we incorporate them into existing polynomial spectral GNNs, such as GPRGNN (Chien et al. 2021), BernNet (He et al. 2021), and JacobiConv (Wang and Zhang 2022).
Insertion into GPR-GNN. GPR-GNN (Chien et al. 2021) directly assigns a learnable coefficient to each order of the normalized adjacency matrix $\mathbf{\hat{A}}$, and its polynomial filter is defined as:
$\mathbf{z}=\sum_{k=0}^{K}\gamma_{k}\mathbf{\hat{A}}^{k}=\mathbf{U}g_{\gamma,K}(\Lambda)\mathbf{U}^{T},$ (6)
where $g_{\gamma,K}(\Lambda)$ is an element-wise operation, and $g_{\gamma,K}(x)=\sum_{k=0}^{K}\gamma_{k}x^{k}$. GPR-GNN represents the simplest form of polynomial spectral GNN, assigning a single scalar coefficient to each propagation step. Therefore, we directly incorporate heterophily basis vectors into GPR-GNN in ascending order of their polynomial order:
$\mathbf{z}=\sum_{k=0}^{K}\gamma_{k}\left(\beta\mathbf{\hat{A}}^{k}\mathbf{x}+(1-\beta)\mathbf{u}_{k}\right),$ (7)
where $\beta$ is a tunable hyperparameter.
Insertion into BernNet. BernNet (He et al. 2021) expresses the filtering operation with Bernstein polynomials and forces all coefficients to be positive, and its filter is defined as:
$\mathbf{z}=\sum_{k=0}^{K}\theta_{k}\frac{1}{2^{K}}\left(\begin{array}[]{c}K\\ k\end{array}\right)(2\mathbf{I}-\mathbf{L})^{K-k}\mathbf{L}^{k}\mathbf{x}.$ (8)
For BernNet, each term is a product involving both $2\mathbf{I}-\mathbf{L}$ and $\mathbf{L}$. We insert the $k$-th heterophily basis vector $\mathbf{u}_{k}$ into the $k$-th order of $\mathbf{L}$:
$\mathbf{z}=\sum_{k=0}^{K}\theta_{k}[\beta\frac{1}{2^{K}}\left(\begin{array}[]{c}K\\ k\end{array}\right)(2\mathbf{I}-\mathbf{L})^{K-k}\mathbf{L}^{k}\mathbf{x}+(1-\beta)\mathbf{u}_{k}].$ (9)
Table 2: Dataset Statistics
Dataset Children History Fitness Cornell Texas Washington Wisconsin Cora Pubmed Citeseer
Nodes 76,875 41,551 173,055 191 187 229 265 2,708 19,717 3,186
Edges 1,554,578 358,574 1,773,500 292 310 394 510 10,556 88,648 8,450
Feat 768 768 768 768 768 768 768 1,433 500 3,703
Classes 24 12 13 5 5 5 5 7 3 6
Homophily 0.42 0.66 0.90 0.12 0.06 0.15 0.17 0.81 0.80 0.78
Insertion into JacobiConv. JacobiConv (Wang and Zhang 2022) proposes a Jacobi basis to adapt a wide range of weight functions due to its orthogonality and flexibility. The iterative process of the Jacobi basis can be defined as:
$P_{0}^{a,b}(x)=1,$ (10)
$P_{1}^{a,b}(x)=0.5a-0.5b+(0.5a+0.5b+1)x,$
$P_{k}^{a,b}(x)=(2k+a+b-1)$
$\cdot\frac{(2k+a+b)(2k+a+b-2)x+a^{2}-b^{2}}{2k(k+a+b)(2k+a+b-2)}P_{k-1}^{a,b}(x)$
$-\frac{(k+a-1)(k+b-1)(2k+a+b)}{k(k+a+b)(2k+a+b-2)}P_{k-2}^{a,b}(x),$
where $a$ and $b$ are tunable hyperparameters. Unlike GPR-GNN and BernNet, JacobiConv adopts an individual filter function for each output dimension $l$:
$\mathbf{Z}_{:l}=\sum_{k=0}^{K}\alpha_{kl}P_{k}^{a,b}(\mathbf{\hat{A}})(\mathbf{X}\mathbf{W})_{:l}.$ (11)
Similarly, we incorporate the corresponding heterophily basis vector $\mathbf{u}_{k}$ into each polynomial order of JacobiConv as follows:
$\mathbf{Z}_{:l}=\sum_{k=0}^{K}\alpha_{kl}[\beta P_{k}^{a,b}(\mathbf{\hat{A}})(\mathbf{X}\mathbf{W})_{:l}+(1-\beta)\mathbf{u}_{k}].$ (12)
4.3 Training Objective
After integrating the LLM-estimated homophily ratio into existing polynomial spectral filters, node classification can be performed using various polynomial-based spectral GNNs. Notably, the proposed method introduces no additional trainable parameters; it simply incorporates heterophily-aware basis vectors guided by the estimated homophily.
We adopt a multi-layer perceptron (MLP) with parameter $\theta$ to predict the label distribution:
$\hat{\mathbf{y}}=\operatorname{MLP}\left(\mathbf{Z};\theta\right),$ (13)
where $\hat{\mathbf{y}}$ is the predicted label distribution. Then, we optimize the cross-entropy loss of the node $j$:
$\mathcal{L}=\sum_{j\in\mathcal{V}_{\text{train }}}\operatorname{CrossEntropy}\left(\hat{\mathbf{y}}^{j},\mathbf{y}^{j}\right),$ (14)
where $\mathcal{V}_{\text{train }}$ is the training node set, and $\mathbf{y}^{j}$ is the ground-truth one-hot label vector of node $j$.
5 Experiment
In this section, to fully evaluate the performance of the proposed model, we present a series of comprehensive experiments to answer the following research questions (RQs):
  • RQ1: Can the performance of polynomial spectral GNNs be improved by incorporating homophily estimated by LLMs?
  • RQ2: How does the homophily estimated by LLMs compare to other estimation methods in enhancing polynomial spectral GNNs?
  • RQ3: Does the LLM-based homophily estimation incur lower inference cost?
  • RQ4: What is the impact of the key hyperparameter $\beta$ on model performance?
5.1 Experimental Setup
Datasets. We select ten graph datasets with text attributes , including three citation networks (Cora, Citeseer, and Pubmed), four webpage networks (Cornell, Texas, Washington, and Wisconsin), and three Amazon co-purchase networks (Children, History, and Fitness). These datasets cover both homophilic and heterophilic graph structures. The statistics of these datasets are summarized in Table 2.
Table 3: Performance (%) on Various Datasets (mean with standard deviation as subscript)
Method Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
MLP 80.85±0.00 84.91±0.00 68.38±1.48 79.82±1.52 54.90±0.04 83.91±0.14 79.57±0.06 79.52±0.32 71.68±1.68 87.43±0.11
GCN 57.45±4.26 35.85±3.77 37.61±10.36 57.02±1.52 56.47±0.10 84.82±0.05 90.86±0.06 86.22±0.57 78.58±0.81 85.02±0.21
GraphSAGE 78.72±2.13 78.62±1.09 69.23±4.44 76.32±2.64 54.39±1.32 84.51±0.19 88.55±2.07 86.65±0.77 76.96±1.36 88.80±0.31
GAT 44.68±11.85 38.36±4.75 30.77±10.26 51.75±15.20 49.78±4.31 82.79±1.18 90.51±0.68 84.93±0.76 78.16±0.45 85.04±0.38
APPNP 68.79±2.45 54.72±0.00 57.26±1.48 62.28±1.52 45.34±0.27 81.42±0.07 80.33±0.36 88.01±0.64 81.40±0.33 87.58±0.03
[…]
UniFilter 88.61±4.16 84.52±3.01 69.00±4.12 79.38±3.65 47.55±0.16 86.99±0.84 87.81±4.74 87.19±1.73 77.57±2.10 85.55±3.63
GPRGNN 82.98±2.13 79.87±4.75 68.38±2.96 81.58±2.63 60.29±0.46 86.03±0.07 92.38±0.07 88.01±0.19 78.16±0.33 89.11±0.19
GPRGNNPLUS 87.23±0.00 81.76±2.18 70.94±2.96 83.33±1.52 61.00±0.20 86.47±0.13 92.58±0.04 88.07±0.21 79.94±0.41 89.53±0.09
BernNet 79.43±1.23 84.91±1.89 67.52±1.48 80.70±1.52 58.76±0.17 84.93±0.30 91.83±0.16 82.41±0.60 72.83±1.48 88.54±0.36
BernNetPLUS 90.78±1.23 85.53±2.18 71.79±5.13 83.33±1.52 60.80±0.19 85.66±0.09 92.44±0.06 87.58±1.06 78.89±1.28 89.37±0.08
JacobiConv 78.01±2.45 84.28±1.09 64.96±3.92 79.82±3.04 59.63±0.46 85.36±0.13 90.70±0.17 84.99±1.01 72.88±1.13 88.04±0.10
JacobiConvPLUS 80.85±0.00 84.91±1.89 68.38±1.48 80.70±1.52 59.67±0.05 85.79±0.07 91.04±0.26 88.25±0.39 78.63±1.48 89.48±0.13
ChebNetII 82.27±1.00 85.53±0.89 68.38±1.21 84.21±2.15 56.74±0.07 84.76±0.10 83.49±0.02 81.01±0.12 72.37±0.24 88.34±0.13
ChebNetIIPLUS 87.94±1.00 85.53±0.89 69.23±2.09 84.21±2.15 58.76±0.12 85.79±0.05 88.22±0.17 83.53±0.17 74.12±0.36 88.34±0.10
Settings. We adopt the experimental setup used in CS-TAG (Yan et al. 2023) and follow its standard data split to ensure a fair comparison. Specifically, the training/validation/test sets are divided as 60%/20%/20% for all datasets except Fitness, for which the split is 20%/10%/70%. We use the accuracy metric as an evaluation indicator. All experiments are performed three times, and we report the average results and their corresponding standard errors. All experiments are conducted on a machine with 3 NVIDIA A5000 24GB GPUs and Intel(R) Xeon(R) Silver 4310 2.10 GHz CPU.
Baselines. To thoroughly evaluate the effectiveness of the proposed method, in addition to the three polynomial spectral GNN backbones—GPR-GNN (Chien et al. 2021), BernNet (He et al. 2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022)—we also include four classical GNN models and a Multi-Layer Perceptron (MLP) as baselines: GCN (Kipf and Welling 2017), GAT (Veličković et al. 2019), GraphSAGE (Hamilton, Ying, and Leskovec 2017), APPNP (Gasteiger, Bojchevski, and Günnemann 2019), TFE-GNN (Duan et al. 2024), and UniFilter (Huang et al. 2024).
5.2 Main Results (RQ1)
Table 3 presents the node classification results across ten datasets. As observed, the proposed method consistently outperforms GPRGNN, BernNet, JacobiConv, and ChebNetII by a significant margin. Specifically, it achieves an average improvement of 1.14% over GPRGNN, 4.51% over BernNet, 2.51% over JacobiConv, and 2.41% over ChebNetII. These results clearly demonstrate the effectiveness of the proposed method. Moreover, although recent methods such as ChebNetII and TFE-GNN have shown competitive performance, our method still surpasses them. This highlights the strong representational power of LLM-enhanced spectral GNNs.
Figure 2: Ablation study of proposed method.
5.3 Ablation Analysis (RQ2)
This subsection aims to evaluate the advantage of using LLM-predicted homophily in polynomial spectral GNNs. We consider the following two alternative variants for comparison:
  1. 1.
    Training a MLP to predict homophily and enhance the spectral GNN.
  2. 2.
    Directly using the homophily computed from the training set to enhance the spectral GNN.
Figure 2 compares the performance of spectral GNNs enhanced with homophily estimated by different methods on four datasets  [^1] Please refer to Appendix B for more results.. As shown, the homophily predicted by LLMs yields the best performance in nearly all cases, with the only exception being BernNet on the Texas dataset. This demonstrates the superiority of LLM-estimated homophily over that derived from MLPs or directly from the training set, likely due to the stronger reasoning capability of LLMs. In addition, MLP-based homophily estimation generally outperforms the one directly obtained from the training set, indicating that learning homophily via a trainable model is more effective than relying on limited labeled data. However, MLP-based variants still fall short of the LLM-based approach, highlighting that stronger models can generate more informative homophily signals, which in turn lead to better performance when integrated into spectral GNNs.
Table 4: Token and cost statistics on each dataset (million tokens / USD)
Type Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
Output 0.12M/0.07\$ 0.13M/0.08\$ 0.14M/0.09\$ 0.13M/0.08\$ 0.17M/0.10\$ 0.16M/0.10\$ 0.15M/0.09\$ 0.17M/0.10\$ 0.16M/0.10\$ 0.18M/0.11\$
Input 0.47M/0.07\$ 0.67M/0.10\$ 0.56M/0.08\$ 0.44M/0.07\$ 0.38M/0.06\$ 0.36M/0.05\$ 0.09M/0.01\$ 0.22M/0.03\$ 0.24M/0.04\$ 0.40M/0.06\$
Total 0.59M/0.14\$ 0.80M/0.18\$ 0.71M/0.17\$ 0.57M/0.14\$ 0.55M/0.16\$ 0.53M/0.15\$ 0.24M/0.10\$ 0.39M/0.13\$ 0.40M/0.13\$ 0.58M/0.17\$
Table 5: Per-epoch training time (ms) and total training time (s) comparisons on various datasets.
Method Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
GPRGNN 13.00/2.60 12.30/2.46 11.80/2.36 13.40/2.67 230.80/46.16 94.90/18.97 472.70/94.54 15.00/3.00 31.40/6.29 32.60/6.52
GPRGNNPLUS 13.60/2.72 14.00/2.81 13.90/2.77 13.90/2.78 274.00/54.80 119.10/23.83 575.40/115.08 18.80/3.76 41.40/8.28 40.60/8.12
BernNet 37.50/7.50 36.90/7.38 38.60/7.72 35.90/7.18 1235.30/247.06 445.70/89.15 2449.40/489.88 50.20/10.04 132.10/26.42 128.30/25.67
BernNetPLUS 43.50/8.69 40.90/8.17 43.80/8.76 43.50/8.71 1303.50/260.70 473.80/94.77 2596.20/519.23 53.70/10.73 143.10/28.63 138.20/27.65
JacobiConv 26.80/5.36 26.80/5.36 26.80/5.36 26.80/5.35 56.80/11.37 33.90/6.77 88.00/17.60 28.80/5.76 28.10/5.63 30.10/6.02
JacobiConvPLUS 28.10/5.63 28.60/5.72 27.90/5.58 28.00/5.59 56.50/11.31 34.40/6.88 86.30/17.26 30.50/6.09 29.80/5.96 31.30/6.26
ChebNetII 49.40/9.88 50.20/10.05 49.60/9.93 49.20/9.83 274.70/54.95 137.00/27.41 574.50/114.89 51.80/10.35 70.40/14.08 68.80/13.76
ChebNetIIPLUS 50.00/10.01 50.70/10.13 50.60/10.12 51.20/10.24 324.20/64.84 164.50/32.90 685.70/137.13 54.80/10.95 79.50/15.89 76.90/15.38
Figure 3: Effect of hyperparameter $\beta$ on model performance.
5.4 Efficiency Studies (RQ3)
To evaluate the efficiency of the proposed method, we measure both its monetary cost and time overhead. Since our method employs low-cost GPT-4o Mini API calls to estimate homophily levels on different datasets, a small monetary expense is incurred. Table 4 reports the input, output, and total monetary cost per dataset. As shown, the cost for each dataset  [^2] Please refer to Appendix C for hyperparameter sensitivity results on additional datasets. remains below \$0.2, indicating that our method achieves high model performance at a minimal financial expense.
To further assess the time efficiency of our approach, we compare the per-epoch and total training time with those of the original polynomial spectral GNNs. As shown in Table 5, the runtime of our method remains similar to the original GNNs across all ten datasets, suggesting that the proposed approach does not increase time complexity.
In practice, the only additional computation introduced by the proposed method, compared to standard spectral GNNs, lies in the construction of heterophily-aware basis functions based on the predicted homophily levels. This step is shared across all graph convolution layers and can thus be precomputed for different polynomial orders. During training, the precomputed basis vectors can be directly loaded to significantly reduce computational overhead. Furthermore, it is worth emphasizing that our approach does not rely on costly fine-tuning or local model deployment, highlighting both the efficiency and practical applicability of the proposed method.
5.5 Parameter Sensitivity Analysis (RQ4)
In the Methodology Section, we introduce a key hyperparameter $\beta$ to balance the contribution between the original polynomial basis and the heterophily-aware basis constructed using predicted homophily. A lower value of $\beta$ indicates a greater reliance on the heterophily-aware basis, while a higher $\beta$ emphasizes the original polynomial basis. Figure 3 illustrates the sensitivity of hyperparameter $\beta$ on four datasets. As shown, different polynomial models exhibit distinct trends depending on the dataset. For example, BernNetPLUS demonstrates an upward trend in the Washington and Texas datasets. This may be attributed to BernNet’s strong capacity to approximate arbitrary filters, making the original polynomial basis more influential. In contrast, GPRGNNPLUS and JacobiConvPLUS appear largely insensitive to changes in $\beta$, suggesting that both the original and heterophily-aware bases are equally effective at capturing meaningful filters. Conversely, on the Cora and Citeseer datasets, all three models show a generally decreasing trend, indicating that for homophilic graphs, incorporating more heterophily-aware basis functions—constructed based on predicted homophily—can lead to improved performance.
6 Conclusion
This paper is the first to investigate the potential of LLMs in enhancing spectral GNNs. We propose a novel framework that utilizes LLM-predicted homophily to construct a heterophily basis, which is then integrated into existing spectral GNN architectures. The approach begins by sampling a subset of edges from the graph and employing an LLM to infer whether connected node pairs belong to the same class, thereby estimating homophily. This estimated homophily guides the construction of a heterophily-aware representation, enabling spectral GNNs to handle both homophilic and heterophilic graphs better.
Extensive experiments on multiple benchmark datasets demonstrate that our method consistently improves the performance of various spectral GNNs, while incurring minimal computational and monetary overhead. These results highlight the promising synergy between LLMs and spectral graph learning, and open up new directions for future research at the intersection of language models and graph neural architectures.

### v2

@@
-Can Large Language Models Improve Spectral Graph Neural Networks?
+Enhancing Spectral Graph Neural Networks with LLM-Predicted Homophily
 Abstract
-Spectral Graph Neural Networks (SGNNs) have attracted significant attention due to their ability to approximate arbitrary filters.
-They typically rely on supervision from downstream tasks to adaptively learn appropriate filters.
-However, under label-scarce conditions, SGNNs may learn suboptimal filters, leading to degraded performance.
-Meanwhile, the remarkable success of Large Language Models (LLMs) has inspired growing interest in exploring their potential within the GNN domain.
-This naturally raises an important question: Can LLMs help overcome the limitations of SGNNs and enhance their performance?
-In this paper, we propose a novel approach that leverages LLMs to estimate the homophily of a given graph.
-The estimated homophily is then used to adaptively guide the design of polynomial spectral filters, thereby improving the expressiveness and adaptability of SGNNs across diverse graph structures.
-Specifically, we introduce a lightweight pipeline in which the LLM generates homophily-aware priors, which are injected into the filter coefficients to better align with the underlying graph topology.
-Extensive experiments on benchmark datasets demonstrate that our LLM-driven SGNN framework consistently outperforms existing baselines under both homophilic and heterophilic settings, with minimal computational and monetary overhead.
+Spectral Graph Neural Networks (SGNNs) have achieved remarkable performance in tasks such as node classification due to their ability to learn flexible filters.
+Typically, these filters are learned under the supervision of downstream tasks, enabling SGNNs to adapt to diverse structural patterns.
+However, in scenarios with limited labeled data, SGNNs often struggle to capture the optimal filter shapes, resulting in degraded performance, especially on graphs with heterophily.
+Meanwhile, the rapid progress of Large Language Models (LLMs) has opened new possibilities for enhancing graph learning without modifying graph structure or requiring task-specific training.
+In this work, we propose a novel framework that leverages LLMs to estimate the homophily level of a graph and uses this global structural prior to guide the construction of spectral filters.
+Specifically, we design a lightweight and plug-and-play pipeline where a small set of labeled node pairs is formatted as natural language prompts for the LLM, which then predicts the graph’s homophily ratio.
+This estimated value informs the spectral filter basis, enabling SGNNs to adapt more effectively to both homophilic and heterophilic structures.
+Extensive experiments on multiple benchmark datasets demonstrate that our LLM-assisted spectral framework consistently improves performance over strong SGNN baselines.
+Importantly, this enhancement incurs negligible computational and monetary cost, making it a practical solution for real-world graph applications.
 1 Introduction
-The past decade has witnessed great success and prosperity of graph neural networks (GNNs) in diverse data science and engineering scenarios, such as traffic network (Hu et al.
-2019; Jiang and Luo 2022), abnormal detection (Tang et al.
-2022, 2024a), relational databases (Cappuzzo, Papotti, and Thirumuruganathan 2020; Huang et al.
-2022) and recommender systems (Sharma et al.
+Graph Neural Networks (GNNs) have demonstrated impressive performance across a wide range of applications, including traffic forecasting (Ahmed et al.
+2024; Jiang and Luo 2022), anomaly detection (Dong et al.
+2025; Tang et al.
+2024a), and recommendation systems (Sharma et al.
 2024; Gao et al.
 2023).
-Existing GNNs can be broadly categorized into spatial GNNs and spectral GNNs.
-Spatial GNNs often adopt a message-passing mechanism to learn node representations by aggregating neighbor node features.
-In contrast, spectral GNNs map node features to a new desired space by selectively attenuating or amplifying the Fourier coefficients induced by the normalized Laplacian matrix.
-This study primarily centers on the realm of spectral GNNs.
-In real-world scenarios, graphs often exhibit varying degrees of homophily.
-In homophilic graphs, nodes of the same class are more likely to be connected, whereas in heterophilic graphs, connections tend to form between nodes of different classes.
-Prior studies [24, 37, 42] have shown that heterophilous structures pose significant challenges for spatial GNNs based on message passing, as these models typically rely on the implicit assumption of homophily.
-To address this issue, many methods have been proposed from the spatial perspective [40].
-In contrast, spectral GNNs [3, 19] with learnable filters offer a promising alternative.
-By learning spectral filters directly from the graph structure, these models can better adapt to heterophilic settings and mitigate the aggregation of noisy information from dissimilar neighbors.
-However, since the labels of real-world graph datasets may be very sparse, existing spectral graph neural network methods will learn inappropriate filters, resulting in suboptimal performance.
-As shown in Figure 1, the optimal filter on the Cora dataset is a low-pass filter due to its high homophily.
-However, classical spectral graph neural networks such as GPRGNN (Chien et al.
-2021), BernNet (He et al.
-2021), and JacobiConv (Wang and Zhang 2022) learn high-pass filters.
-Therefore, the performance of spectral graph neural networks may be suboptimal.
-In recent years, large language models (LLMs), exemplified by GPT-4 (Achiam et al.
-2023), have shown impressive proficiency in comprehending and reasoning over textual information, thereby revolutionizing a wide range of domains, including natural language processing (Zhao et al.
-2023), computer vision (Wu et al.
-2023), and graph representation learning (Ren et al.
-2024)].
-In this work, we explore the potential of language models in spectral graph neural networks.
-To the best of our knowledge, this is the first study to leverage large language models to improve spectral graph neural networks.
-Specifically, we aim to address the following research questions:
-1.
-Can large language models (LLMs) enhance spectral GNNs?
-LLMs are well known for their prowess in understanding natural language, while spectral GNNs excel at node classification in heterophilic graphs.
-Effectively combining the text comprehension capabilities of LLMs with the filtering power of spectral GNNs presents a significant challenge.
-Existing efforts to integrate GNNs and LLMs typically take one of two approaches: either feeding textual graphs directly into LLMs—limiting the model’s ability to leverage graph structure, or fine-tuning LLMs for graph-specific tasks, which introduces substantial computational and deployment costs.
-2.
-Can LLMs effectively guide and be integrated into existing spectral GNNs?
-Spectral GNNs have the powerful ability to approximate arbitrary filters, which are particularly suitable for both homophilic and heterophilic graphs.
-However, these models usually rely on supervision from downstream tasks to learn appropriate spectral filters, which can lead to suboptimal filter learning in label-scarce or weakly-supervised settings.
-On the other hand, LLMs demonstrate strong performance in language understanding and text-driven tasks, as seen in models like LLM-GNN (Chen et al.
-2023b).
-This opens up a promising and novel direction: using LLM-generated predictions to assist and improve spectral GNNs.
-In this paper, we propose an LLM-based enhancement framework for spectral GNNs.
-Specifically, to address the first research question: spectral GNNs may learn inappropriate filters due to sparse labels, while the homophily ratio serves as a global characteristic of a graph dataset.
-We naturally leverage LLMs to predict the homophily ratio, enabling spectral GNNs to perceive global graph properties even under limited supervision.
-Notably, predicting homophily only requires sampling around 100 edges, making the cost per dataset as low as \$1, without the need for costly fine-tuning or deployment.
-For the second research question, we incorporate the predicted homophily ratio into existing polynomial spectral GNNs to enhance their performance.
-Specifically, we construct various heterophily-aware bases according to the predicted homophily, and then integrate them into polynomial spectral GNNs.
-This integration allows the model to fully utilize homophily information, boosting overall performance.
+By encoding both structural and attribute information into low-dimensional representations, GNNs enable effective learning on non-Euclidean data.
+Existing architectures can be broadly classified into spatial and spectral models.
+Spatial GNNs propagate messages between neighboring nodes and typically assume homophily—i.e., connected nodes share similar labels.
+This assumption, however, often breaks down in heterophilic graphs, where neighbors may hold dissimilar labels, leading to performance degradation (Ma et al.
+2022; Xu et al.
+2022).
+In contrast, spectral GNNs operate in the frequency domain, applying polynomial filters to modulate graph signals based on the spectrum of the Laplacian matrix.
+This design enables fine-grained control over low- and high-frequency information, offering improved adaptability to both homophilic and heterophilic structures (Chien et al.
+2021; He et al.
+2021; Wang and Zhang 2022).
+In particular, models with learnable filters have shown strong potential in capturing complex structural patterns.
+However, recent studies (Xu et al.
+2024; Huang et al.
+2024) have pointed out that the effectiveness of spectral GNNs heavily depends on the quality and shape of their learned filters, which are typically supervised by downstream node classification tasks.
+In practice, acquiring sufficient labeled nodes is both time-consuming and expensive, especially in large-scale or sensitive domains such as biology or cybersecurity.
+As a result, models trained with limited labels often struggle to learn appropriate filters, leading to suboptimal generalization.
+A promising yet underexplored direction is to incorporate external prior knowledge or auxiliary estimation to guide spectral filter learning.
+In this context, large language models (LLMs), such as GPT-4 (Achiam et al.
+2023), emerge as powerful candidates.
+These models have exhibited remarkable capabilities in semantic understanding, reasoning, and generalization across multiple domains, including natural language (Zhao et al.
+2023), vision (Wu et al.
+2023), and even structured data such as graphs (Ren et al.
+2024; Jin et al.
+2024).
+Their ability to perform few-shot or zero-shot inference makes them particularly appealing for graph tasks with limited supervision.
+While prior work has attempted to combine LLMs and GNNs, existing approaches largely fall into two categories: (i) LLM-centric, where graph information is encoded into text and processed entirely by LLMs (Tang et al.
+2024b; Zhang et al.
+2024), often ignoring structural nuances; and (ii) GNN-centric, where LLMs are used to augment node features or generate edges (He et al.
+2024; Xie et al.
+2023), which requires fine-tuning or task-specific training.
+These approaches either sacrifice structural expressiveness or incur considerable computational overhead, limiting their scalability and generality.
+In this work, we propose a new perspective: using LLMs not to directly process graph data, but to assist spectral GNNs by estimating a key global structural property—graph homophily.
+Our intuition is that LLMs, given a small set of sampled edge-label pairs in natural language format, can effectively predict the overall homophily level of the graph due to their strong generalization and semantic reasoning abilities.
+The predicted homophily ratio can then guide the construction of spectral filters, improving GNN adaptability in both homophilic and heterophilic settings.
+To this end, we propose a lightweight and general framework that leverages LLMs to estimate the homophily ratio of a graph.
+Specifically, we query the LLM using a small set of sampled node pairs and their labels, and aggregate its predictions to approximate the graph-level homophily ratio.
+This approach is cost-efficient (less than \$0.2 per dataset), requires no expensive model fine-tuning, and can be used with any black-box LLM.
+We then construct heterophily-aware polynomial bases guided by the predicted homophily ratio and integrate them into existing spectral GNNs.
+This allows the model to adapt its filter shape to the underlying graph structure.
 Our contributions are summarized as follows:
-  • To the best of our knowledge, this is the first work to investigate the integration of LLMs into spectral GNNs.
-We show that LLMs can effectively predict the homophily ratio to guide spectral GNNs, thereby boosting their performance without the need for fine-tuning or additional deployment overhead.
-  • We introduce a novel LLM-based enhancement framework for spectral GNNs, which is compatible with various LLMs and polynomial spectral architectures.
-By leveraging the global semantic understanding of LLMs, our method significantly improves the adaptability and performance of spectral GNNs across diverse scenarios.
-  • Extensive experiments on multiple benchmark datasets demonstrate that our method consistently improves the performance of existing spectral GNNs, while incurring minimal computational and financial overhead.
+  • Novel Integration of LLMs into Spectral GNNs.
+We are the first to explore the integration of LLMs into spectral GNNs, using LLMs to estimate graph homophily and guide spectral filtering, without requiring fine-tuning, retraining, or graph structure modification.
+  • A General and Lightweight Enhancement Framework.
+We propose a plug-and-play framework compatible with various spectral GNN architectures and off-the-shelf LLMs.
+Our method leverages the semantic reasoning abilities of LLMs to inject global structural priors into spectral models, enhancing their adaptability to different homophily regimes.
+  • Extensive Empirical Validation.
+We conduct comprehensive experiments on multiple benchmark datasets, demonstrating that our framework consistently improves the performance of strong spectral GNN baselines, while incurring minimal computational and monetary overhead.
 2 Related Work
@@
 Spectral GNNs.
-According to whether the filter can be learned, the spectral GNNs can be divided into pre-defined filters and learnable filters.
-In the category of pre-defined filters, GCN (Kipf and Welling 2017) uses a simplified first-order Chebyshev polynomial.
-APPNP (Gasteiger, Bojchevski, and Günnemann 2019) utilizes Personalized Page Rank to set the weight of the filter.
-In the category of learnable filters.
-ChebNet (Defferrard, Bresson, and Vandergheynst 2016) uses Chebyshev polynomials with learnable coefficients.
-GPR-GNN (Chien et al.
-2021) extends APPNP by directly parameterizing its weights.
-BernNet (He et al.
-2021) uses Bernstein polynomials to learn filters and forces all coefficients positive.
-JacobiConv (Wang and Zhang 2022) adopts an orthogonal and flexible Jacobi basis to accommodate a wide range of weight functions.
-ChebNetII (He, Wei, and Wen 2022) uses Chebyshev interpolation to learn filters.
-Specformer (Bo et al.
-2023) performs self-attention in the spectral domain to learn a set-to-set spectral filter.
-Recently, some works have applied spectral GNNs to node-level filtering.
-For example, DSF (Guo et al.
-2023) proposes a novel diversified spectral filtering framework that automatically learns node-specific filter weights.
-UniFilter (Huang et al.
-2024) integrates heterophilic and homophilic bases to construct a universal polynomial basis, UniBasis, which partially alleviates the problems of over-smoothing and over-squashing.
+Spectral GNNs can be broadly categorized based on their filtering strategies.
+Early methods like GCN (Kipf and Welling 2017) and APPNP (Gasteiger, Bojchevski, and Günnemann 2019) use pre-defined filters based on fixed polynomial formulations.
+Later approaches introduce learnable filters for greater flexibility: ChebNet (Defferrard, Bresson, and Vandergheynst 2016), GPR-GNN (Chien et al.
+2021), and BernNet (He et al.
+2021) adopt various polynomial bases with trainable coefficients; JacobiConv (Wang and Zhang 2022) leverages orthogonal Jacobi polynomials; ChebNetII (He, Wei, and Wen 2022) uses Chebyshev interpolation; and Specformer (Bo et al.
+2023) introduces spectral self-attention.
+Recently, UniFilter (Huang et al.
+2024) unifies heterophilic and homophilic bases to construct a universal polynomial filter that alleviates over-smoothing and over-squashing.
 LLM for Graph.
 Existing research on applying large language models (LLMs) to graph learning can be broadly categorized into GNN-centric and LLM-centric approaches.
-GNN-centric methods leverage LLMs to extract node features from raw data, and then use GNNs for downstream prediction tasks (He et al.
+GNN-centric methods typically leverage LLMs to extract node features from raw data, and then use GNNs for downstream prediction tasks (He et al.
 2024; Xie et al.
@@
 2024b; Zhang et al.
-2024a).
-Some studies also employ LLMs to assign edge weights in text-attributed graphs (Sun et al.
-2023; Ling et al.
 2024).
-However, these methods are not specifically designed for heterophilic graphs.
+However, these methods are based on the homophily assumption.
 Beyond these, a few studies aim to enhance GNNs by leveraging LLMs to identify meaningful or noisy edges.
@@
 2024) utilizes LLMs for graph structure learning by detecting and removing noisy connections; LLM4RGNN (Zhang et al.
-2024b) identifies malicious and critical edges to improve the adversarial robustness of GNNs; and LLM4HeG (Wu et al.
-2024) integrates LLMs into GNNs for heterophilic TAGs through edge discrimination and adaptive edge reweighting.
-In contrast to existing methods, the proposed method uniquely focuses on leveraging LLMs to improve and enhance spectral GNNs, which has been underexplored in prior work.
+2025) identifies malicious and critical edges to improve the adversarial robustness of GNNs; and LLM4HeG (Wu et al.
+2025) integrates LLMs into GNNs for heterophilic graphs through edge discrimination and adaptive edge reweighting.
+Existing approaches typically tailor LLM usage to specific graph tasks or graph structural modifications.
+In contrast, our work introduces a general and lightweight framework that utilizes LLMs to estimate the homophily ratio for enhancing spectral GNNs, without requiring LLM fine-tuning or changes to the graph.
+This represents a novel integration pathway between LLMs and spectral graph learning
 3 Preliminaries
 3.1 Spectral GNN
-Assume we are given an undirected homogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{n}\right\}$ denotes the set of $n$ nodes, $\mathcal{E}$ is the edge set, and $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the node feature matrix.
+Assume we are given an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{n}\right\}$ denotes the set of $n$ nodes, $\mathcal{E}$ is the edge set, and $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the node feature matrix.
 The corresponding adjacency matrix is $\mathbf{A}\in\{0,1\}^{n\times n}$, where $\mathbf{A}_{ij}=1$ if there is an edge between nodes $v_{i}$ and $v_{j}$, and $\mathbf{A}_{ij}=0$ otherwise.
@@
 Let $\mathbf{\hat{L}}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ denote the eigen-decomposition of $\mathbf{\hat{L}}$, where $\mathbf{U}$ is the matrix of eigenvectors and $\boldsymbol{\Lambda}=\text{diag}([\lambda_{1},\lambda_{2},\ldots,\lambda_{n}])$ is the diagonal matrix of eigenvalues.
-Spectral GNNs are based on the Fourier transform in signal processing.
+Spectral GNN is based on the Fourier transform in signal processing.
 The Fourier transform of a graph signal $\mathbf{x}$ is given by $\hat{\mathbf{x}}=\mathbf{U}^{\top}\mathbf{x}$, and its inverse is $\mathbf{x}=\mathbf{U}\hat{\mathbf{x}}$.
@@
 where $K$ is the order of the polynomial $h(\cdot)$ and $h_{k}$ is the coefficient of the $k$-th order term. Thus, Eq. (1) can be rewritten as:
-$\mathbf{Z}=\mathbf{H}\mathbf{X}=h(\mathbf{\hat{L}})\mathbf{X}=\mathbf{U}h(\boldsymbol{\Lambda})\mathbf{U}^{\top}\mathbf{X},$ (3)
+$\mathbf{Z}$ $=\mathbf{H}\mathbf{X}=h(\mathbf{\hat{L}})\mathbf{X}=\mathbf{U}h(\boldsymbol{\Lambda})\mathbf{U}^{\top}\mathbf{X}$ (3)
+$=\mathbf{U}\left(\begin{array}[]{ccc}h(\lambda_{1})&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&h(\lambda_{n})\end{array}\right)\mathbf{U}^{\top}\mathbf{X},$
 where $\mathbf{Z}$ is the output (prediction) matrix.
-According to Eq.
-(2), Eq.
-(3), and recent studies (Chen et al.
-2023a), a $K$-th order polynomial in spectral GNNs is equivalent to aggregating information from $K$-hop neighbors in spatial GNNs.
 3.2 Homophily
 The homophily metric measures the degree of association between connected nodes.
-Homophily can be measured in various ways, such as node-level homophily, edge-level homophily, and class-level homophily.
-The widely adopted edge homophily (Zhu et al.
-2020) is defined as follows:
+Homophily can be measured in various ways, such as node-level homophily (Pei et al.
+2020), edge-level homophily (Zhu et al.
+2020), and class-level homophily (Lim et al.
+2021).
+The widely adopted edge homophily is defined as follows:
 $\mathcal{H}_{\text{edge }}(\mathcal{G})=\frac{1}{|\mathcal{E}|}\sum_{(u,v)\in\mathcal{E}}\mathbf{1}\left(y_{u}=y_{v}\right),$ (4)
 where $\mathbf{1}(\cdot)$ is the indicator function, i.e., $\mathbf{1}(\cdot)=1$ if the condition holds, otherwise $\mathbf{1}(\cdot)=0$.
-$y_{u}$ is the label of node $u$, and $y_{v}$ is the label of node $v$ $|\mathcal{E}|$ is the size of the edge set.
+$y_{u}$ and $y_{v}$ denote the labels of nodes $u$ and $v$, respectively.
+$|\mathcal{E}|$ is the size of the edge set.
 4 Methodology
 Figure 1: The overall framework of the proposed model.
-First, the proposed method samples a subset of edges (node pairs) from the dataset.
+First, we sample a subset of edges (node pairs) from the dataset.
 Then we use LLM to predict whether the connected nodes belong to the same class.
@@
 Finally, a heterophily basis is constructed using the predicted homophily and integrated into existing polynomial spectral GNNs to perform the node classification task.
-This section describes the proposed method, which consists of two modules: (1) using an LLM to predict the homophily, and (2) incorporating the predicted homophily into existing spectral GNNs.
+This section describes the proposed method, which consists of two modules: (1) using an LLM to predict the homophily ratio, and (2) incorporating the predicted homophily into existing spectral GNNs.
 The following provides a detailed introduction to these two modules.
 4.1 Estimating Homophily Using LLM
-It is well established that, for a graph, the homophily ratio measures the degree of correlation among neighboring nodes and reflects the proportion of edges connecting nodes of the same class.
-Consequently, accurately estimating the homophily ratio is critical for better leveraging the graph’s structural properties and enabling targeted spectral filtering.
-For example, in a highly homophilic graph, a polynomial spectral GNN should ideally learn a low-pass filter.
-Table 1: Comparison of homophily predicted by the LLM with varying prompts, with indicators for whether CoT and Most-voting are used.
-Prompt Type Cora Citeseer Pubmed History
-COT Most-voting
-× × 0.41 0.63 0.72 0.58
-× 0.51 0.76 0.67 0.65
-× 0.37 0.63 0.81 0.55
-0.70 0.81 0.83 0.77
-Ground Truth 0.81 0.78 0.80 0.66
-For simplicity, we adopt the widely used edge-level homophily as our estimation target, as Eq.
-(4).
-Specifically, we aim to estimate the proportion of homophily edges (i.e., edges linking nodes of the same class) among all edges.
-However, it is impractical to evaluate the homophily of every edge, especially in large-scale graphs with millions of edges, due to the high inference cost associated with LLMs.
-Therefore, we propose to estimate the overall homophily ratio by sampling a subset of edges.
-To achieve accurate estimation of the homophily ratio, it is important to assess the reliability of predictions, such as through calibrated confidence scores.
+It is well established that the homophily ratio of a graph quantifies the degree of label similarity among neighboring nodes, reflecting the proportion of edges that connect nodes belonging to the same class.
+Consequently, accurate estimation of this ratio is essential for leveraging structural properties of the graph and enabling effective spectral filtering.
+For instance, in highly homophilic graphs, a polynomial spectral GNN is expected to learn a low-pass filter (Bo et al.
+2021).
+Table 1: Comparison of homophily predicted by the LLM with varying prompts, with indicators for whether CoT and Most-voting (Most) are used.
+The percentage reflects the closeness to ground truth (higher is better).
+Prompt Type Cora Citeseer Pubmed Average
+CoT Most
+× × 0.41(49%) 0.63(81%) 0.72(90%) 73%
+× 0.51(63%) 0.76(97%) 0.67(84%) 81%
+× 0.37(46%) 0.63(81%) 0.81(99%) 75%
+0.70(86%) 0.81(96%) 0.83(98%) 93%
+Ground Truth 0.81 0.78 0.80
+For simplicity, we adopt the widely used edge-level homophily (Zhu et al.
+2020) as the target to estimate.
+Specifically, we aim to determine the proportion of homophilic edges—edges connecting nodes of the same class—among all edges.
+However, exhaustively evaluating every edge is impractical, especially for large-scale graphs with millions of edges, due to the high inference cost of LLMs.
+Thus, we estimate the overall homophily by sampling a representative subset of edges.
+To ensure accurate estimation of homophily, it is important to improve the quality of individual edge predictions produced by the LLM.
 Motivated by recent advances (Chen et al.
-2023b) in generating calibrated confidence from LLMs, we investigate the following strategies:
-  • Vanilla prompting: Directly querying the model for its confidence.
-  • Reasoning-based prompting(CoT): Guiding the model through annotation generation using techniques such as chain-of-thought and multi-step reasoning.
-  • Consistency-based prompting (Most-voting): Querying the model multiple times and selecting the most frequent prediction via majority voting.
-  • Hybrid prompting: Combining Reasoning-based and consistency-based methods to enhance robustness.
-As shown in Table 1, both reasoning-based prompting and consistency-based prompting contribute to improved performance in predicting homophily.
-Therefore, we adopt a hybrid prompting strategy that integrates their respective strengths to further enhance the accuracy of homophily estimation.
-For example, the prompt used for the Cora dataset is designed as follows, while prompts for other datasets can be found in Appendix A.
+2024) in enhancing prediction reliability via prompt design, we explore the following strategies:
+  • Vanilla prompting: Directly querying the model without any additional reasoning or consistency mechanisms.
+  • Reasoning-based prompting (CoT): Encouraging the model to perform step-by-step reasoning using chain-of-thought prompts.
+  • Consistency-based prompting (Most-voting): Querying the model multiple times and using majority voting to determine the final prediction.
+  • Hybrid prompting: Combining both reasoning and consistency to improve robustness and accuracy.
+As shown in Table 1, both reasoning and consistency-based prompting enhance homophily prediction accuracy.
+Therefore, we adopt a hybrid prompting strategy to combine their complementary strengths.
+The following is an example prompt:
+System: You are a chatbot expert in text classification.
+User: We have two node texts from the following categories: [categories list].
+The texts are as follows: Node $v_{i}\rightarrow${Title,Abstract}.
+Node $v_{j}\rightarrow${Title,Abstract}.
+Please tell me whether they belong to the same category or not after reasoning step by step.
+After obtaining initial predictions using reasoning-based prompting, we further enhance reliability via consistency-based voting.
+Specifically, each sampled edge is queried five times, and the final label is determined via majority voting.
+Let $y_{(u,v)}\in\{0,1\}$ denote the final label for edge $(u,v)$, where 1 indicates nodes $u$ and $v$ belong to the same class (i.e.
+homophilic edge), and 0 otherwise.
+The decision rule is defined as:
+$y_{(u,v)}=\left\{\begin{array}[]{ll}1&\text{if }r_{(u,v)}\geq 3,\\ 0&\text{if }r_{(u,v)}<3,\end{array}\right$.
+where $r_{(u,v)}$ denotes the number of responses labeling edge $(u,v)$ as homophilic among the five queries. Finally, the edge-level homophily $\hat{h}$ is estimated as:
+$\hat{h}=\frac{|\{(u,v)\in\mathcal{E}_{s}\mid y_{(u,v)}=1\}|}{|\mathcal{E}_{s}|},$ (5)
+where $\mathcal{E}_{s}$ denotes the set of all sampled edges.
+4.2 Incorporating Homophily into Polynomial Spectral Filters
+After estimating the homophily ratio using an LLM, we integrate the predicted homophily into a polynomial spectral GNN framework.
+To enable spectral GNNs to exploit the homophily information inferred in Eq.
+(5), we construct a set of heterophily-aware basis vectors, inspired by UniFilter (Huang et al.
+2024).
+Specifically, the angle between each pair of basis vectors is defined as:
 […]
-After obtaining the homophily labels for a subset of edges using reasoning-based prompting, we further apply consistency-based prompting by querying each sample five times.
-If at least three out of five predictions yield the same answer, we consider it the final prediction for that sample.
-The process is illustrated as follows:
-$y_{e}=\left\{\begin{array}[]{ll}1&\text{ if }r_{e}\geq 3\\ 0&\text{ if }r_{e}<3\end{array},\right.$
-Once the predictions for each sampled edge are obtained, we compute the predicted edge-level homophily $\hat{h}$ as follows:
-$\hat{h}=\frac{|\{(u,v)\in\mathcal{E}_{s}\mid\text{LLM}(u,v)=\text{True}\}|}{|\mathcal{E}_{s}|},$ (5)
-where $\mathcal{E}_{s}$ denotes the set of all sampled edges.
-4.2 Combination of Homophily with Existing Polynomial Filters
-After predicting the homophily ratio using an LLM, we incorporate the estimated value into an existing polynomial spectral GNN.
-To enable spectral GNNs to leverage the homophily information predicted by Eq.
-(5), we construct a set of heterophily-aware basis vectors inspired by UniFilter (Huang et al.
-2024).
-Specifically, the angle between each pair of basis vectors is defined as follows:
-$\theta=\frac{\pi}{2}(1-\hat{h}).$
-Thus, the angle between different heterophily-aware basis vectors $\mathbf{u}_{i}$ and $\mathbf{u}_{j}$ is given by:
+[…]
 $\mathbf{u}_{i}\cdot\mathbf{u}_{j}=\begin{cases}\cos\theta=\cos\left(\frac{(1-\hat{h})\pi}{2}\right)&\text{ if }i\neq j,\\ 1&\text{ if }i=j.\end{cases}$
@@
 2021), BernNet (He et al.
-2021), and JacobiConv (Wang and Zhang 2022).
+2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022).
+The details are as follows:
 Insertion into GPR-GNN.
@@
 where $\beta$ is a tunable hyperparameter.
+Table 2: Dataset Statistics
+Dataset Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
+Nodes 229 191 187 76,875 41,551 173,055 2,708 3,186 19,717
+Edges 394 292 310 1,554,578 358,574 1,773,500 5,429 4,277 44,335
+Feat 768 768 768 768 768 768 1,433 3,703 500
+Classes 5 5 5 24 12 13 7 6 3
+Homophily 0.15 0.12 0.06 0.42 0.66 0.90 0.81 0.78 0.80
 Insertion into BernNet. BernNet (He et al. 2021) expresses the filtering operation with Bernstein polynomials and forces all coefficients to be positive, and its filter is defined as:
@@
 $\mathbf{z}=\sum_{k=0}^{K}\theta_{k}[\beta\frac{1}{2^{K}}\left(\begin{array}[]{c}K\\ k\end{array}\right)(2\mathbf{I}-\mathbf{L})^{K-k}\mathbf{L}^{k}\mathbf{x}+(1-\beta)\mathbf{u}_{k}].$ (9)
-Table 2: Dataset Statistics
-Dataset Children History Fitness Cornell Texas Washington Wisconsin Cora Pubmed Citeseer
-Nodes 76,875 41,551 173,055 191 187 229 265 2,708 19,717 3,186
-Edges 1,554,578 358,574 1,773,500 292 310 394 510 10,556 88,648 8,450
-Feat 768 768 768 768 768 768 768 1,433 500 3,703
-Classes 24 12 13 5 5 5 5 7 3 6
-Homophily 0.42 0.66 0.90 0.12 0.06 0.15 0.17 0.81 0.80 0.78
 Insertion into JacobiConv.
@@
 $P_{0}^{a,b}(x)=1,$ (10)
-$P_{1}^{a,b}(x)=0.5a-0.5b+(0.5a+0.5b+1)x,$
+[…]
 $P_{k}^{a,b}(x)=(2k+a+b-1)$
@@
 $\mathbf{Z}_{:l}=\sum_{k=0}^{K}\alpha_{kl}[\beta P_{k}^{a,b}(\mathbf{\hat{A}})(\mathbf{X}\mathbf{W})_{:l}+(1-\beta)\mathbf{u}_{k}].$ (12)
+Insertion Details of ChebNetII.
+To mitigate overfitting and reduce the Runge phenomenon, ChebNetII (He, Wei, and Wen 2022) enhances the original Chebyshev (Defferrard, Bresson, and Vandergheynst 2016) polynomial approximation by reparameterizing the filter response as a learnable coefficient $\gamma_{j}$ , as shown below:
+$\mathbf{z}=\frac{2}{K+1}\sum_{k=0}^{K}\sum_{j=0}^{K}\gamma_{j}T_{k}\left(x_{j}\right)T_{k}(\mathbf{L})\mathbf{x},$ (13)
+where $x_{j}=\cos((j+1/2)\pi/(K+1))$ are the Chebyshev nodes of $T_{K+1}$. The recursive definition of the Chebyshev polynomial $T_{k}(x)$ is given as follows:
+$T_{0}(x)$ $=1,$ (14)
+$T_{1}(x)$ $=x,$
+$T_{k}(x)$ $=2xT_{k-1}(x)-T_{k-2}(x).$
+We incorporate the heterophily basis vector $\mathbf{u}_{k}$ into the corresponding $k$-th term of the ChebNetII, resulting in:
+$\mathbf{z}=\frac{2}{K+1}\sum_{k=0}^{K}\left[\beta\sum_{j=0}^{K}\gamma_{j}T_{k}\left(x_{j}\right)T_{k}(\mathbf{L})\mathbf{x}+(1-\beta)\mathbf{u}_{k}\right],$ (15)
 4.3 Training Objective
 After integrating the LLM-estimated homophily ratio into existing polynomial spectral filters, node classification can be performed using various polynomial-based spectral GNNs.
-Notably, the proposed method introduces no additional trainable parameters; it simply incorporates heterophily-aware basis vectors guided by the estimated homophily.
+Notably, the proposed method introduces no additional trainable parameters.
 We adopt a multi-layer perceptron (MLP) with parameter $\theta$ to predict the label distribution:
@@
 $\mathcal{L}=\sum_{j\in\mathcal{V}_{\text{train }}}\operatorname{CrossEntropy}\left(\hat{\mathbf{y}}^{j},\mathbf{y}^{j}\right),$ (17)
-where $\mathcal{V}_{\text{train }}$ is the training node set, and $\mathbf{y}^{j}$ is the ground-truth one-hot label vector of node $j$.
+[…]
 5 Experiment
 In this section, to fully evaluate the performance of the proposed model, we present a series of comprehensive experiments to answer the following research questions (RQs):
-  • RQ1: Can the performance of polynomial spectral GNNs be improved by incorporating homophily estimated by LLMs?
-  • RQ2: How does the homophily estimated by LLMs compare to other estimation methods in enhancing polynomial spectral GNNs?
-  • RQ3: Does the LLM-based homophily estimation incur lower inference cost?
-  • RQ4: What is the impact of the key hyperparameter $\beta$ on model performance?
+  • RQ1: How does the proposed method perform compared to its corresponding spectral GNN counterparts?
+  • RQ2: How effective is LLM-predicted homophily compared to that estimated by MLPs or directly derived from the training set?
+  • RQ3: Does the LLM-based homophily estimation incur negligible time and money cost?
+  • RQ4: How does the key hyperparameter $\beta$ influence the model’s performance?
 5.1 Experimental Setup
 Datasets.
-We select ten graph datasets with text attributes , including three citation networks (Cora, Citeseer, and Pubmed), four webpage networks (Cornell, Texas, Washington, and Wisconsin), and three Amazon co-purchase networks (Children, History, and Fitness).
-These datasets cover both homophilic and heterophilic graph structures.
+We select nine graph datasets with text attributes (Yan et al.
+2023; Guo et al.
+2024) , including three citation networks (Cora, Citeseer, and Pubmed), three webpage networks (Cornell, Texas, and Washington), and three Amazon co-purchase networks (Children, History, and Fitness).
 The statistics of these datasets are summarized in Table 2.
-Table 3: Performance (%) on Various Datasets (mean with standard deviation as subscript)
-Method Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
-MLP 80.85±0.00 84.91±0.00 68.38±1.48 79.82±1.52 54.90±0.04 83.91±0.14 79.57±0.06 79.52±0.32 71.68±1.68 87.43±0.11
-GCN 57.45±4.26 35.85±3.77 37.61±10.36 57.02±1.52 56.47±0.10 84.82±0.05 90.86±0.06 86.22±0.57 78.58±0.81 85.02±0.21
-GraphSAGE 78.72±2.13 78.62±1.09 69.23±4.44 76.32±2.64 54.39±1.32 84.51±0.19 88.55±2.07 86.65±0.77 76.96±1.36 88.80±0.31
-GAT 44.68±11.85 38.36±4.75 30.77±10.26 51.75±15.20 49.78±4.31 82.79±1.18 90.51±0.68 84.93±0.76 78.16±0.45 85.04±0.38
-APPNP 68.79±2.45 54.72±0.00 57.26±1.48 62.28±1.52 45.34±0.27 81.42±0.07 80.33±0.36 88.01±0.64 81.40±0.33 87.58±0.03
-[…]
-UniFilter 88.61±4.16 84.52±3.01 69.00±4.12 79.38±3.65 47.55±0.16 86.99±0.84 87.81±4.74 87.19±1.73 77.57±2.10 85.55±3.63
-GPRGNN 82.98±2.13 79.87±4.75 68.38±2.96 81.58±2.63 60.29±0.46 86.03±0.07 92.38±0.07 88.01±0.19 78.16±0.33 89.11±0.19
-GPRGNNPLUS 87.23±0.00 81.76±2.18 70.94±2.96 83.33±1.52 61.00±0.20 86.47±0.13 92.58±0.04 88.07±0.21 79.94±0.41 89.53±0.09
-BernNet 79.43±1.23 84.91±1.89 67.52±1.48 80.70±1.52 58.76±0.17 84.93±0.30 91.83±0.16 82.41±0.60 72.83±1.48 88.54±0.36
-BernNetPLUS 90.78±1.23 85.53±2.18 71.79±5.13 83.33±1.52 60.80±0.19 85.66±0.09 92.44±0.06 87.58±1.06 78.89±1.28 89.37±0.08
-JacobiConv 78.01±2.45 84.28±1.09 64.96±3.92 79.82±3.04 59.63±0.46 85.36±0.13 90.70±0.17 84.99±1.01 72.88±1.13 88.04±0.10
-JacobiConvPLUS 80.85±0.00 84.91±1.89 68.38±1.48 80.70±1.52 59.67±0.05 85.79±0.07 91.04±0.26 88.25±0.39 78.63±1.48 89.48±0.13
-ChebNetII 82.27±1.00 85.53±0.89 68.38±1.21 84.21±2.15 56.74±0.07 84.76±0.10 83.49±0.02 81.01±0.12 72.37±0.24 88.34±0.13
-ChebNetIIPLUS 87.94±1.00 85.53±0.89 69.23±2.09 84.21±2.15 58.76±0.12 85.79±0.05 88.22±0.17 83.53±0.17 74.12±0.36 88.34±0.10
+Table 3: Performance (%) on Various Datasets (mean accuracy with standard deviation as subscript)
+Method Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
+MLP 80.85±0.00 68.38±1.48 79.82±1.52 54.90±0.04 83.91±0.14 79.57±0.06 79.52±0.32 71.68±1.68 87.43±0.11
+GCN 57.45±4.26 37.61±10.36 57.02±1.52 56.47±0.10 84.82±0.05 90.86±0.06 86.22±0.57 78.58±0.81 85.02±0.21
+GraphSAGE 78.72±2.13 69.23±4.44 76.32±2.64 54.39±1.32 84.51±0.19 88.55±2.07 86.65±0.77 76.96±1.36 88.80±0.31
+GAT 44.68±11.85 30.77±10.26 51.75±15.20 49.78±4.31 82.79±1.18 90.51±0.68 84.93±0.76 78.16±0.45 85.04±0.38
+APPNP 68.79±2.45 57.26±1.48 62.28±1.52 45.34±0.27 81.42±0.07 80.33±0.36 88.01±0.64 81.40±0.33 87.58±0.03
+TFEGNN 68.09±9.19 66.67±5.54 64.04±1.24 45.44±0.83 81.19±0.03 83.31±1.02 85.50±0.27 77.83±0.52 88.25±0.17
+UniFilter 88.61±4.16 69.00±4.12 79.38±3.65 47.55±0.16 86.99±0.84 87.81±4.74 87.19±1.73 77.57±2.10 85.55±3.63
+GPRGNN 82.98±2.13 68.38±2.96 81.58±2.63 60.29±0.46 86.03±0.07 92.38±0.07 88.01±0.19 78.16±0.33 89.11±0.19
+GPRGNNPLUS 87.23±0.00 70.94±2.96 83.33±1.52 61.00±0.20 86.47±0.13 92.58±0.04 88.07±0.21 79.94±0.41 89.53±0.09
+BernNet 79.43±1.23 67.52±1.48 80.70±1.52 58.76±0.17 84.93±0.30 91.83±0.16 82.41±0.60 72.83±1.48 88.54±0.36
+BernNetPLUS 90.78±1.23 71.79±5.13 83.33±1.52 60.80±0.19 85.66±0.09 92.44±0.06 87.58±1.06 78.89±1.28 89.37±0.08
+JacobiConv 78.01±2.45 64.96±3.92 79.82±3.04 59.63±0.46 85.36±0.13 90.70±0.17 84.99±1.01 72.88±1.13 88.04±0.10
+JacobiConvPLUS 80.85±0.00 68.38±1.48 80.70±1.52 59.67±0.05 85.79±0.07 91.04±0.26 88.25±0.39 78.63±1.48 89.48±0.13
+ChebNetII 82.27±1.00 68.38±1.21 84.21±2.15 56.74±0.07 84.76±0.10 83.49±0.02 81.01±0.12 72.37±0.24 88.34±0.13
+ChebNetIIPLUS 87.94±1.00 69.23±2.09 84.21±2.15 58.76±0.12 85.79±0.05 88.22±0.17 83.53±0.17 74.12±0.36 88.34±0.10
 Settings.
@@
 All experiments are conducted on a machine with 3 NVIDIA A5000 24GB GPUs and Intel(R) Xeon(R) Silver 4310 2.10 GHz CPU.
+The LLM used for generating homophily ratio predictions is GPT-4o-mini (Hurst et al.
+2024).
 Baselines.
-To thoroughly evaluate the effectiveness of the proposed method, in addition to the three polynomial spectral GNN backbones—GPR-GNN (Chien et al.
+To thoroughly evaluate the effectiveness of the proposed method, in addition to the four polynomial spectral GNN backbones—GPR-GNN (Chien et al.
 2021), BernNet (He et al.
-2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022)—we also include four classical GNN models and a Multi-Layer Perceptron (MLP) as baselines: GCN (Kipf and Welling 2017), GAT (Veličković et al.
+2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022)—we also include six classical GNN models and a Multi-Layer Perceptron (MLP) as baselines: GCN (Kipf and Welling 2017), GAT (Veličković et al.
 2019), GraphSAGE (Hamilton, Ying, and Leskovec 2017), APPNP (Gasteiger, Bojchevski, and Günnemann 2019), TFE-GNN (Duan et al.
@@
 2024).
+For more experimental details, including comprehensive descriptions of the baselines, datasets, and hyperparameter settings, please refer to Appendix A.
 5.2 Main Results (RQ1)
-Table 3 presents the node classification results across ten datasets.
-As observed, the proposed method consistently outperforms GPRGNN, BernNet, JacobiConv, and ChebNetII by a significant margin.
-Specifically, it achieves an average improvement of 1.14% over GPRGNN, 4.51% over BernNet, 2.51% over JacobiConv, and 2.41% over ChebNetII.
-These results clearly demonstrate the effectiveness of the proposed method.
-Moreover, although recent methods such as ChebNetII and TFE-GNN have shown competitive performance, our method still surpasses them.
-This highlights the strong representational power of LLM-enhanced spectral GNNs.
+Table 3 reports the node classification accuracies on nine benchmark datasets.
+As observed, our LLM-enhanced variants (denoted with the PLUS suffix) consistently outperform their original counterparts, including GPRGNN, BernNet, JacobiConv, and ChebNetII, across all datasets.
+Concretely, the average performance gains over the original models are substantial: +1.17% for GPRGNN, +4.93% for BernNet, +2.70% for JacobiConv, and +2.68% for ChebNetII.
+These consistent improvements clearly demonstrate the effectiveness of incorporating LLM-predicted homophily into spectral GNNs.
+In addition, despite the competitive performance of recent spectral methods such as UniFilter (Huang et al.
+2024) and TFE-GNN (Duan et al.
+2024), our approach still achieves superior results across diverse datasets.
+Importantly, our method introduces no graph topology modifications or task-specific training, yet enhances the adaptability of spectral GNNs to different structural patterns, including both homophilic and heterophilic graphs.
+Overall, the results demonstrate that incorporating global structural priors via LLMs can significantly boost the representational power of spectral GNNs, offering a simple yet effective enhancement to existing models.
 Figure 2: Ablation study of proposed method.
 5.3 Ablation Analysis (RQ2)
-This subsection aims to evaluate the advantage of using LLM-predicted homophily in polynomial spectral GNNs. We consider the following two alternative variants for comparison:
-  1. 1.
-    Training a MLP to predict homophily and enhance the spectral GNN.
-  2. 2.
-    Directly using the homophily computed from the training set to enhance the spectral GNN.
-Figure 2 compares the performance of spectral GNNs enhanced with homophily estimated by different methods on four datasets  [^1] Please refer to Appendix B for more results..
-As shown, the homophily predicted by LLMs yields the best performance in nearly all cases, with the only exception being BernNet on the Texas dataset.
-This demonstrates the superiority of LLM-estimated homophily over that derived from MLPs or directly from the training set, likely due to the stronger reasoning capability of LLMs.
-In addition, MLP-based homophily estimation generally outperforms the one directly obtained from the training set, indicating that learning homophily via a trainable model is more effective than relying on limited labeled data.
-However, MLP-based variants still fall short of the LLM-based approach, highlighting that stronger models can generate more informative homophily signals, which in turn lead to better performance when integrated into spectral GNNs.
+We evaluate the impact of incorporating LLM-predicted homophily into spectral GNNs (LLM-based) by comparing it with two variants: (1) MLP-based, which trains an MLP to classify each edge as homophilic or heterophilic for estimating overall homophily; (2) Train-label-based, which directly computes homophily from the proportion of homophilic edges among labeled node pairs in the training set.
+Figure 2 presents the performance of these variants on four datasets.
+As shown, LLM-predicted homophily consistently achieves the best results, with the only exception being BernNet on the Texas dataset.
+This underscores the advantage of LLMs in generating more reliable structural priors, likely due to their stronger reasoning capabilities.
+Notably, MLP-based estimation also outperforms the direct use of training-set homophily, indicating that learning homophily signals is more effective than relying on limited labeled data.
+However, it still lags behind the LLM-based variant, highlighting that the quality of the estimated homophily plays a crucial role in guiding spectral filter construction.
 Table 4: Token and cost statistics on each dataset (million tokens / USD)
-Type Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
-Output 0.12M/0.07\$ 0.13M/0.08\$ 0.14M/0.09\$ 0.13M/0.08\$ 0.17M/0.10\$ 0.16M/0.10\$ 0.15M/0.09\$ 0.17M/0.10\$ 0.16M/0.10\$ 0.18M/0.11\$
-Input 0.47M/0.07\$ 0.67M/0.10\$ 0.56M/0.08\$ 0.44M/0.07\$ 0.38M/0.06\$ 0.36M/0.05\$ 0.09M/0.01\$ 0.22M/0.03\$ 0.24M/0.04\$ 0.40M/0.06\$
-Total 0.59M/0.14\$ 0.80M/0.18\$ 0.71M/0.17\$ 0.57M/0.14\$ 0.55M/0.16\$ 0.53M/0.15\$ 0.24M/0.10\$ 0.39M/0.13\$ 0.40M/0.13\$ 0.58M/0.17\$
+Type Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
+Input 0.47/0.07\$ 0.56/0.08\$ 0.44/0.07\$ 0.38/0.06\$ 0.36/0.05\$ 0.09/0.01\$ 0.22/0.03\$ 0.24/0.04\$ 0.40/0.06\$
+Output 0.12/0.07\$ 0.14/0.09\$ 0.13/0.08\$ 0.17/0.10\$ 0.16/0.10\$ 0.15/0.09\$ 0.17/0.10\$ 0.16/0.10\$ 0.18/0.11\$
+Total 0.59/0.14\$ 0.71/0.17\$ 0.57/0.14\$ 0.55/0.16\$ 0.53/0.15\$ 0.24/0.10\$ 0.39/0.13\$ 0.40/0.13\$ 0.58/0.17\$
 Table 5: Per-epoch training time (ms) and total training time (s) comparisons on various datasets.
-Method Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
-GPRGNN 13.00/2.60 12.30/2.46 11.80/2.36 13.40/2.67 230.80/46.16 94.90/18.97 472.70/94.54 15.00/3.00 31.40/6.29 32.60/6.52
-GPRGNNPLUS 13.60/2.72 14.00/2.81 13.90/2.77 13.90/2.78 274.00/54.80 119.10/23.83 575.40/115.08 18.80/3.76 41.40/8.28 40.60/8.12
-BernNet 37.50/7.50 36.90/7.38 38.60/7.72 35.90/7.18 1235.30/247.06 445.70/89.15 2449.40/489.88 50.20/10.04 132.10/26.42 128.30/25.67
-BernNetPLUS 43.50/8.69 40.90/8.17 43.80/8.76 43.50/8.71 1303.50/260.70 473.80/94.77 2596.20/519.23 53.70/10.73 143.10/28.63 138.20/27.65
-JacobiConv 26.80/5.36 26.80/5.36 26.80/5.36 26.80/5.35 56.80/11.37 33.90/6.77 88.00/17.60 28.80/5.76 28.10/5.63 30.10/6.02
-JacobiConvPLUS 28.10/5.63 28.60/5.72 27.90/5.58 28.00/5.59 56.50/11.31 34.40/6.88 86.30/17.26 30.50/6.09 29.80/5.96 31.30/6.26
-ChebNetII 49.40/9.88 50.20/10.05 49.60/9.93 49.20/9.83 274.70/54.95 137.00/27.41 574.50/114.89 51.80/10.35 70.40/14.08 68.80/13.76
-ChebNetIIPLUS 50.00/10.01 50.70/10.13 50.60/10.12 51.20/10.24 324.20/64.84 164.50/32.90 685.70/137.13 54.80/10.95 79.50/15.89 76.90/15.38
+Method Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
+GPRGNN 13.0/2.6 11.8/2.4 13.4/2.7 230.8/46.2 94.9/19.0 472.7/94.5 15.0/3.0 31.4/6.3 32.6/6.5
+GPRGNNPLUS 13.6/2.7 13.9/2.8 13.9/2.8 274.0/54.8 119.1/23.8 575.4/115.1 18.8/3.8 41.4/8.3 40.6/8.1
+BernNet 37.5/7.5 38.6/7.7 35.9/7.2 1235.3/247.1 445.7/89.2 2449.4/489.9 50.2/10.0 132.1/26.4 128.3/25.7
+BernNetPLUS 43.5/8.7 43.8/8.8 43.5/8.7 1303.5/260.7 473.8/94.8 2596.2/519.2 53.7/10.7 143.1/28.6 138.2/27.7
+JacobiConv 26.8/5.4 26.8/5.4 26.8/5.4 56.8/11.4 33.9/6.8 88.0/17.6 28.8/5.8 28.1/5.6 30.1/6.0
+JacobiConvPLUS 28.1/5.6 27.9/5.6 28.0/5.6 56.5/11.3 34.4/6.9 86.3/17.3 30.5/6.1 29.8/6.0 31.3/6.3
+ChebNetII 49.4/9.9 49.6/9.9 49.2/9.8 274.7/55.0 137.0/27.4 574.5/114.9 51.8/10.4 70.4/14.1 68.8/13.8
+ChebNetIIPLUS 50.0/10.0 50.6/10.1 51.2/10.2 324.2/64.8 164.5/32.9 685.7/137.1 54.8/11.0 79.5/15.9 76.9/15.4
 Figure 3: Effect of hyperparameter $\beta$ on model performance.
 5.4 Efficiency Studies (RQ3)
-To evaluate the efficiency of the proposed method, we measure both its monetary cost and time overhead.
-Since our method employs low-cost GPT-4o Mini API calls to estimate homophily levels on different datasets, a small monetary expense is incurred.
-Table 4 reports the input, output, and total monetary cost per dataset.
-As shown, the cost for each dataset  [^2] Please refer to Appendix C for hyperparameter sensitivity results on additional datasets.
-remains below \$0.2, indicating that our method achieves high model performance at a minimal financial expense.
-To further assess the time efficiency of our approach, we compare the per-epoch and total training time with those of the original polynomial spectral GNNs.
-As shown in Table 5, the runtime of our method remains similar to the original GNNs across all ten datasets, suggesting that the proposed approach does not increase time complexity.
-In practice, the only additional computation introduced by the proposed method, compared to standard spectral GNNs, lies in the construction of heterophily-aware basis functions based on the predicted homophily levels.
-This step is shared across all graph convolution layers and can thus be precomputed for different polynomial orders.
-During training, the precomputed basis vectors can be directly loaded to significantly reduce computational overhead.
-Furthermore, it is worth emphasizing that our approach does not rely on costly fine-tuning or local model deployment, highlighting both the efficiency and practical applicability of the proposed method.
+We assess the efficiency of our method in terms of both monetary and time costs.
+Leveraging low-cost GPT-4o-mini API calls for homophily estimation, the financial overhead remains negligible—under \$0.2 per dataset as shown in Table 4.
+For time efficiency, Table 5 compares our method with baseline spectral GNNs in per-epoch and total training time, showing no significant runtime increase.
+The only extra cost stems from constructing heterophily-aware basis functions, which are precomputed once per polynomial order and reused across layers.
+During training, these basis vectors are simply loaded, adding minimal overhead.
+Notably, our method requires no fine-tuning or local model deployment, making it lightweight and practical.
 5.5 Parameter Sensitivity Analysis (RQ4)
-In the Methodology Section, we introduce a key hyperparameter $\beta$ to balance the contribution between the original polynomial basis and the heterophily-aware basis constructed using predicted homophily.
-A lower value of $\beta$ indicates a greater reliance on the heterophily-aware basis, while a higher $\beta$ emphasizes the original polynomial basis.
-Figure 3 illustrates the sensitivity of hyperparameter $\beta$ on four datasets.
-As shown, different polynomial models exhibit distinct trends depending on the dataset.
-For example, BernNetPLUS demonstrates an upward trend in the Washington and Texas datasets.
-This may be attributed to BernNet’s strong capacity to approximate arbitrary filters, making the original polynomial basis more influential.
-In contrast, GPRGNNPLUS and JacobiConvPLUS appear largely insensitive to changes in $\beta$, suggesting that both the original and heterophily-aware bases are equally effective at capturing meaningful filters.
-Conversely, on the Cora and Citeseer datasets, all three models show a generally decreasing trend, indicating that for homophilic graphs, incorporating more heterophily-aware basis functions—constructed based on predicted homophily—can lead to improved performance.
+Our method introduces a key hyperparameter $\beta$ to balance the original polynomial basis and the heterophily-aware basis informed by predicted homophily.
+Smaller $\beta$ values emphasize the heterophily-aware component, while larger values favor the original basis.
+Figure 3 shows the performance sensitivity to $\beta$ across four representative datasets.
+The results reveal diverse trends.
+For example, in Washington and Texas, BernNetPLUS shows an upward trend, indicating its strong filter capacity benefits from the original basis.
+In contrast, the other three methods maintain stable performance in Washington and Texas, suggesting robustness to $\beta$ and complementary contributions from both bases.
+On datasets like Cora and Citeseer, all models perform better with smaller $\beta$, implying that incorporating heterophily-aware basis functions—guided by LLM-estimated homophily—enhances spectral expressiveness even in largely homophilic graphs.
 6 Conclusion
 This paper is the first to investigate the potential of LLMs in enhancing spectral GNNs.
-We propose a novel framework that utilizes LLM-predicted homophily to construct a heterophily basis, which is then integrated into existing spectral GNN architectures.
-The approach begins by sampling a subset of edges from the graph and employing an LLM to infer whether connected node pairs belong to the same class, thereby estimating homophily.
-This estimated homophily guides the construction of a heterophily-aware representation, enabling spectral GNNs to handle both homophilic and heterophilic graphs better.
+Specifically, we propose a novel method that utilizes LLM-predicted homophily ratio to construct a heterophily basis, which is then integrated into existing spectral GNN architectures, enabling spectral GNNs to handle both homophilic and heterophilic graphs better.
 Extensive experiments on multiple benchmark datasets demonstrate that our method consistently improves the performance of various spectral GNNs, while incurring minimal computational and monetary overhead.
-These results highlight the promising synergy between LLMs and spectral graph learning, and open up new directions for future research at the intersection of language models and graph neural architectures.
+Besides, these results highlight the promising synergy between LLMs and spectral graph learning, and open up new directions for future research at the intersection of larger language models and graph neural networks.
+Appendix A More Experimental Details
+A.1 Datasets
+  • Cora, Citeseer, and Pubmed (Guo et al.
+2024) are three widely used citation network datasets in graph learning.
+In each dataset, nodes represent papers and edges indicate citation links between them.
+Node features are bag-of-words representations of the papers, and labels correspond to document categories.
+  • Cornell, Texas, and Wisconsin (Yan et al.
+2023) are webpage networks collected from computer science departments of different universities.
+In these datasets, nodes represent web pages and edges denote hyperlinks between them.
+We extract node features by encoding the raw webpage content using the pre-trained bert-base-uncased model.
+  • Children and History (Yan et al.
+2023) are two subsets of the Amazon-Books dataset, containing books labeled with the second-level categories “Children” and “History”, respectively.
+Nodes represent books, and edges indicate frequent co-purchase or co-view relationships.
+We use the book titles and descriptions as textual attributes, encoded with the pre-trained roberta-base model.
+  • Fitness (Yan et al.
+2023) is a subset of the Amazon-Sports dataset, containing items labeled with the second-level category “Fitness”.
+Nodes represent fitness-related products, and edges indicate frequent co-purchase or co-view relationships.
+We use the product titles and descriptions as textual features, encoded with the pre-trained roberta-base model.
+A.2 Baselines
+To facilitate understanding of the models used in our experiments, we briefly introduce the backbone polynomial spectral GNN and the baseline GNN methods evaluated in this paper:
+  • GCN (Kipf and Welling 2017) introduces a spectral-based convolution operation on graphs, aggregating information from immediate neighbors with a normalized adjacency matrix.
+  • GraphSAGE (Hamilton, Ying, and Leskovec 2017) samples and aggregates features from a node’s local neighborhood using functions like mean or LSTM, enabling inductive learning on large-scale graphs.
+  • GAT (Veličković et al. 2019) leverages self-attention mechanisms to assign different importance weights to neighbors, enhancing the flexibility of node representation learning.
+  • APPNP (Gasteiger, Bojchevski, and Günnemann 2019) decouples feature transformation and propagation by applying personalized PageRank-based diffusion on top of a neural predictor.
+  • GPRGNN (Chien et al. 2021) extends APPNP by directly assigning a learnable coefficient to each order of the normalized adjacency matrix.
+  • BernNet (He et al. 2021) expresses the filtering operation with Bernstein polynomials and forces all coefficients to be positive.
+  • JacobiConv (Wang and Zhang 2022) proposes a Jacobi basis to adapt a wide range of weight functions due to its orthogonality and flexibility.
+  • ChebNetII (He, Wei, and Wen 2022) addresses overfitting and the Runge phenomenon by enhancing the original Chebyshev polynomial approximation, reparameterizing the filter response as a learnable coefficient.
+  • TFE-GNN (Duan et al. 2024) introduces a triple filter ensemble to combine low-pass, high-pass, and initial features, enabling adaptive extraction of homophily and heterophily.
+  • UniFilter (Huang et al. 2024) unifies heterophilic and homophilic bases to construct a universal polynomial filter that alleviates over-smoothing and over-squashing.
+A.3 Hyper-parameter Settings
+In all experiments, the number of training epochs is set to 1000, and the learning rate is fixed at 0.0001.
+A three-layer MLP with a hidden dimension of 256 is used for feature transformation.
+For each dataset, 100 edges are sampled.
+The key hyperparameter $\beta$ is tuned over the range $\{0.0,0.1,0.2,\dots,1.0\}$.
+The optimal $\beta$ values for each method on each dataset are shown in Table 1.
+Table 6: Optimal $\beta$ values for each method on each dataset.
+Dataset Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
+GPRGNNPLUS 0.5 1.0 1.0 1.0 1.0 0.1 0.7 0.6 0.3
+BernNetPLUS 0.9 0.5 1.0 0.2 0.1 0.2 0.8 0.3 0.5
+JacobiConvPLUS 0.4 0.0 0.0 0.8 0.7 0.6 0.1 0.1 0.1
+ChebnetIIPLUS 0.4 0.8 1.0 0.1 0.1 0.0 0.2 0.8 1.0